OpenStudy (anonymous):

Which property of equality justifies the statement If -4 + k = -3 then k = 1? Answer Addition Property of Equality Multiplication Property of Equality Symmetric Property of Equality Transitive Property of Equality

OpenStudy (anonymous):

-4+k=-3 k=-3-4 k=-7

OpenStudy (anonymous):

k=1 sorry k=-3+4 k=1